# Perceptual compression

### v1

@@ Abstract
Image codecs are typically optimized to trade-off bitrate vs. distortion metrics. At low bitrates, this leads to compression artefacts which are easily perceptible, even when training with perceptual or adversarial losses. To improve image quality, and to make it less dependent on the bitrate, we propose to decode with iterative diffusion models, instead of feed-forward decoders trained using MSE or LPIPS distortions used in most neural codecs. In addition to conditioning the model on a vector-quantized image representation, we also condition on a global textual image description to provide additional context. We dub our model PerCo for “perceptual compression”, and compare it to state-of-the-art codecs at rates from 0.1 down to 0.003 bits per pixel. The latter rate is an order of magnitude smaller than those considered in most prior work. At this bitrate a $512\!\times\!768$ Kodak image is encoded in less than 153 bytes. Despite this ultra-low bitrate, our approach maintains the ability to reconstruct realistic images. We find that our model leads to reconstructions with state-of-the-art visual quality as measured by FID and KID, and that the visual quality is less dependent on the bitrate than previous methods.
@@ 1 Introduction
Traditional image and video codecs are optimized for the rate-distortion function (Shannon, 1948), which minimizes the expected size of the data under a distortion constraint such as mean-squared error. Recent research has developed neural image compression methods that surpass handcrafted image compression codecs in terms of rate-distortion performance (Ballé et al., 2017; 2018; Cheng et al., 2020; He et al., 2022). However, optimization for the rate-distortion function comes at a cost of “realism”, where realism is mathematically related to the statistical fidelity or $f$-divergence between the compressed image distribution and the true image distribution (Blau & Michaeli, 2018; 2019). The typical qualitative manifestation of unrealistic images is blurring.
@@
Generative modeling compensates for artefacts such as blurring by introducing a divergence term (Agustsson et al., 2019; Mentzer et al., 2020; Muckley et al., 2023), typically in the form of an adversarial discriminator loss, which improves human-perceptual performance (Mentzer et al., 2020). Such codecs are called “generative compression” codecs, and are evaluated in terms of the rate-distortion-realism tradeoff (Blau & Michaeli, 2019). A further result of rate-distortion-realism theory is the theoretical possibility of a perfect realism codec, i.e., a codec with no $f$-divergence, and zero FID, across all rates with no more than twofold increase in distortion from the rate-distortion optimal codec. This result motivates research on perfect realism codecs, where original and reconstruction are different from each other, but it is not possible to tell which is which. Such codecs are particularly interesting for extremely low bitrate settings, where existing codecs introduce severe artefacts that are easily perceptible, see Fig. 1. So far, the primary effort towards building such a codec was the work of Theis et al. (2022), who demonstrated lower FID scores than HiFiC (Mentzer et al., 2020) via a combination of a pretrained variational diffusion model and reverse channel coding. However, computational constraints prohibited application to images larger than $64\!\times\!64$ pixels.
@@
In this paper we make further progress towards perfect-realism codecs. Similar to Theis et al. (2022), we leverage a pretrained text-to-image diffusion model; however, rather than using the diffusion model for compression directly, we use it as a decoder in a VQ-VAE-like autoencoder. The (approximate) log-likelihood loss of diffusion models offers an alternative to the use of distortions such as MSE or LPIPS used to train most neural compression models, that make strong assumptions on the conditonal distribution on images given the latents which are inappropriate at low bitrates, and lead to compression artefacts. For decoding, we sample from the conditional diffusion model, which allows us to obtain a set of reconstructions that reflect the uncertainty about the original source image. We augment the quantized image representation with a textual image description that captures the high-level semantic information of the image, and demonstrate that this high-level information is helpful for compression performance at extremely low rates, as low as 0.003 bits per pixel, that are an order of magnitude smaller than those considered in most prior work. Our work is closely related to that of Lei et al. (2023), who also explored text-conditioned diffusion decoders for image compression, and to the best of our knowledge is the only prior work that considers similarly low bitrates. In contrast, their work uses (automatically generated) binary contour maps as a spatial conditioning signal which carry little detail on colors and textures, and leads to reconstructions that are not faithful to the original. Experimentally we observe significantly improved FID and KID scores compared to competing methods. Moreover, we find that FID and KID are much more stable across bitrates than other methods, which aligns with our goal of image compression with perfect realism.
@@ 2 Related work
Neural image compression codecs. Ballé et al. (2017) demonstrated that an end-to-end neural image compression codec could outperform the classical JPEG codec in terms of rate-distortion performance. Ballé et al. (2018) enhance this approach by conditioning the latent on a “hyperprior” that encodes high-level information in the image, greatly improving performance. Follow-up works improved the latent conditioning mechanism (Minnen et al., 2018; Minnen & Singh, 2020; Cheng et al., 2020; He et al., 2022). Optimization for rate-distortion alone leads to distributional mismatch between the compressed image distribution and the natural image distribution (Blau & Michaeli, 2018; 2019). Agustsson et al. (2019) and Mentzer et al. (2020) added an adversarial loss term and demonstrated its usefulness for human perception. Follow-up works have improved the discriminator architecture (Muckley et al., 2023), decode-time realism guidance (Agustsson et al., 2023), or investigated alternative autoencoders with discrete entropy models (El-Nouby et al., 2023).
@@
Diffusion models. Ho et al. (2020) improved the original diffusion probabilistic model (Sohl-Dickstein et al., 2015) that gave it impressive performance on image generation benchmarks. Dhariwal & Nichol (2021) demonstrated superior performance to GANs for image generation with classifier guidance, while Kingma et al. (2021) improved diffusion model likelihood estimation. Conditioning diffusion models on large text-encoder networks enables these models to generate realistic images from natural language prompts (Nichol et al., 2022; Saharia et al., 2022), and defining the diffusion process in an autoencoder latent space reduces the computational cost of the diffusion process (Rombach et al., 2022). In our work we build upon large-scale pretrained text-conditioned diffusion models to benefit from the image generation capabilities embedded in them.
@@
Diffusion generative compressors. Ho et al. (2020) considered compression with diffusion models and reverse channel coding and the corresponding rate-distortion performance, but reported only experiments on $32\!\times\!32$ CIFAR-10 images and MSE distortion. Theis et al. (2022) demonstrated that the reverse channel coding approach gave FID numbers superior to HiFiC for $64\!\times\!64$ ImageNet images. Saharia et al. (2022) introduced a super-resolution diffusion model that removed artefacts from JPEG images. Ghouse et al. (2023) developed a complete version of a residual diffusion codec with images of high quality for a suite of codecs. Hoogeboom et al. (2023) showed similar results using a neural autoencoder baseline codec, and using a separate diffusion model that does not require any additional bits to refine the reconstruction. Yang & Mandt (2023) developed an alternative approach, where a hyperprior-based neural encoder was trained jointly with a diffusion decoder, also showing improved performance to HiFiC. Lei et al. (2023) leverage text-conditioned diffusion models for image compression at very low bitrates, using prompt inversion to encode the image into text, and adding spatial detail via compressed binary contour maps. Similar to Lei et al. (2023), we leverage a text-conditioned diffusion model, but rather than using binary contour maps which carry little appearance information, we condition the diffusion model on more versatile end-to-end learned vector-quantized image features. Moreover, rather than relying on computationally costly iterative prompt inversion (Wen et al., 2023), we rely on a feed-forward image captioning model (Li et al., 2023) to obtain the conditioning text.
@@ 3.1 Overall framework
where $P_{\bm{x}}$ is the data distribution and $P_{{\bm{z}}|{\bm{x}}}$ is the posterior distribution of the quantized codes. The rate term $\mathcal{L}_{\textrm{R}}({\bm{z}})$ estimates the bitrate either by means of scalar quantization and a continuous entropy model (Ballé et al., 2017), or using vector-quantization in combination with discrete entropy models (El-Nouby et al., 2023). For simplicity, in PerCo, we employ vector-quantization combined with a uniform entropy model leading to a constant rate which is determined by the quantization codebook size and the number of elements that are quantized. Therefore, the rate term $\mathcal{L}_{R}$ can be dropped from the optimization formulation. Next, we introduce a new formulation of the distortion term, $\mathcal{L}_{\textrm{D}}(\cdot,\cdot)$, enabling the integration of a pre-trained diffusion model, thereby achieving a higher level of perceptual quality. Our aim with this formulation is to leverage a large pre-trained diffusion model as a robust image prior, enabling realistic image reconstruction even at extremely low bitrates.
@@
Figure 2: Overview of PerCo: the LDM encoder maps an RGB image into the latent space of the diffusion model. The “hyper encoder” then maps the image to a hyper-latent with smaller spatial resolution, which is then vector-quantized and represented as a bitstream using uniform coding. The image captioning model generates a textual description of the input image, which is losslessly compressed, and processed by a text encoder to condition the diffusion model. The diffusion model reconstructs the input image in its latent space conditioned on the output of the text encoder and the hyper encoder. Finally, the LDM decoder maps the latent reconstruction back to RGB pixel space.
@@
Diffusion models, as introduced by Ho et al. (2020), approximate a data distribution by gradually denoising a unit Gaussian random variable. This is achieved by training a denoising function to reverse a diffusion process, which takes an initial sample, denoted as ${\bm{x}}_{0}$, from the data distribution and adds small Gaussian noise over a large number of time steps. In practice, a noise estimator $\epsilon_{\theta}({\bm{x}}_{t},t,{\bm{z}})$ implemented as a neural network is trained to denoise input data points at timesteps $t\in[0,T]$, where ${\bm{x}}_{t}\!=\!\sqrt{\alpha_{t}}{\bm{x}}_{0}\!+\!\sqrt{1-\alpha_{t}}\epsilon$ is a corrupted version of the original data point ${\bm{x}}_{0}$ with Gaussian noise $\epsilon$. The scalar $\alpha_{t}$ fixes the noise strength, varying from $\alpha_{0}\!\simeq\!1$ for no noise to $\alpha_{T}\!\simeq\!0$ for pure Gaussian noise. Finally, ${\bm{z}}$ is an optional condition, such as a text-prompt in text-to-image models. In our case, ${\bm{z}}$ is the quantized representation of the input image. Once the noise estimator is trained, the model can generate samples by drawing Gaussian noise samples, denoted as ${\bm{x}}_{T}\sim\mathcal{N}(0,I)$, and applying iterative denoising for $t\!=\!T,\dots,1$. In practice, to reduce the computational cost, we use a latent diffusion model (LDM) (Rombach et al., 2022), and the variables ${\bm{x}}_{t}$ are defined in the latent space over which the diffusion process is defined.
@@
$\mathcal{L}_{\textrm{Diff}}^{t}\propto\mathbb{E}_{P_{\bm{x}}}\mathbb{E}_{P_{{\bm{z}},{\bm{x}}_{t}|{\bm{x}}}}\mathop{\mathbb{E}}_{\epsilon\sim\mathcal{N}(0,1)}\lVert\epsilon-\epsilon_{\theta}({\bm{x}}_{t},{\bm{z}},t)\rVert_{2}^{2}.$ (3)
We follow Ho et al. (2022), and obtain our final diffusion loss by taking the expectation of Eq. 3 over the steps $t$, ignoring the multiplicative constants.
We use their ${\bm{z}}$ prediction objective, which is more stable during training. The noise estimator does not directly predict the Gaussian noise $\epsilon$, but it predicts ${\bm{v}}_{t}$
such that we have $\epsilon_{\theta}=\sqrt{1-\alpha_{t}}{\bm{x}}_{t}+\sqrt{\alpha_{t}}{\bm{v}}_{t}$.
From the output of the noise estimator, we can estimate of the original data $\hat{{\bm{x}}}_{0}=\sqrt{\alpha_{t}}{\bm{x}}_{t}-\sqrt{1-\alpha_{t}}{\bm{v}}_{t}$. Then, our diffusion loss in Eq. 3 can be augmented with any image-level loss such as LPIPS (Zhang et al., 2018) to improve perceptual image similarity. Our experiments show that this additional term is beneficial at high bitrates.
@@ 3.2 Encoding local and global context
Local spatial encoding. We encode the input image into a quantized tensor proceeding in three stages. First, we employ the LDM encoder $E(\cdot)$ to achieve a first dimension reduction. Assuming an input image of resolution, $512\!\times\!512$, we obtain a feature map of dimension $4\!\times\!64\!\times\!64$. Second, we add a lightweight “hyper encoder” $H$ composed of several convolutional layers to project the LDM features to a lower spatial resolution tensor ${\bm{H}}_{s}$. The spatial resolution, $h\!\times\!w$, of ${\bm{H}}_{s}$ is chosen according to the target bitrate. Third, we proceed to quantization of ${\bm{H}}_{s}$ to obtain ${\bm{z}}_{g}$ via vector-quatization. As in VQ-VAE (van den Oord et al., 2017; Razavi et al., 2019), each vector ${\bm{h}}_{s}$ in ${\bm{H}}_{s}$ —one for every spatial location— is mapped to an element in a codebook learned end-to-end via minimization of the vector quantization loss:
@@
where $sg(.)$ is the stop-gradient operation, and ${\bm{z}}_{q}$ is the mapping of ${\bm{h}}_{s}$ to its closest codebook entry.
To achieve higher compression performance, we use several techniques that were shown to be successful in guaranteeing better codebook usage by Yu et al. (2022). First, we lower the codebook dimension to 32 and project down the encoder values to this lower dimension. We also found it beneficial to replace the codebook loss by exponential moving average on the codebooks and thus the first term in Eq. 4 is no longer used.
The quantized encoder output is then upsampled to $64\!\times\!64$ resolution, to be channel-wise concatenated to the input of the noise estimation network of the diffusion model. The hyper-network is trained end-to-end while finetuning the diffusion model.
For simplicity, we opt for a uniform coding scheme to map the quantized latents to a bitstream, using $\log_{2}V$ bits to encode each element in the $h\!\times\!w$ sized hyper latent encoding, where $V$ is the codebook size. Thus resulting in a bitstream of length $hw\log_{2}V$ bits.
@@
Global encoding with image captioning. While ${\bm{z}}_{l}$ can accurately encode local information, our experiments show that relying on ${\bm{z}}_{l}$ alone leads to unsatisfying realism. To address this issue we also transmit a text-based encoding of the image ${\bm{z}}_{g}$ in order to provide a global description of the image. The second motivation for using text is that its ASCII representation can be efficiently transmitted. For this purpose, we use an off-the-shelf state-of-the-art image captioning model, such as BLIP-2 (Li et al., 2023) or IDEFICS (Laurançon et al., 2023), which we keep fixed when training our model. Note that user-provided captions can also be employed, as further discussed in our experiments. Similar to Lei et al. (2023), we losslessly compress the image caption using Lempel-Ziv coding as implemented in the zlib library (zli, ) to obtain ${\bm{z}}_{g}$.
@@
Decoding with a diffusion model. The quantized local image features in ${\bm{z}}_{l}$ are fed to the denoising U-Net of the diffusion model via concatenation with the latent features ${\bm{x}}_{t}$ of the current time step. To adapt to this change of dimension, we extend the kernel of the first convolutional layer of the U-Net by adding a number of channels that corresponds to the channel dimension of ${\bm{z}}_{l}$ which is randomly initialized. The global encoding ${\bm{z}}_{g}$ is losslessly decoded and passed to the diffusion model through cross-attention layers of the pre-trained diffusion model following Rombach et al. (2022).
Furthermore, we found it beneficial to use classifier-free guidance (Ho & Salimans, 2021), which we apply at inference time for the text conditioning ${\bm{z}}_{g}$, i.e. we contrast conditioning on local featues ${\bm{z}}_{l}$ alone vs. conditioning on both ${\bm{z}}_{l}$ and ${\bm{z}}_{g}$. This leads to the noise estimate:
$\hat{\epsilon}_{\theta}=\epsilon_{\theta}({\bm{x}}_{t},({\bm{z}}_{l},\emptyset),t)+\lambda_{s}~{}\big{(}\epsilon_{\theta}({\bm{x}}_{t},({\bm{z}}_{l},{\bm{z}}_{g}),t)-\epsilon_{\theta}({\bm{x}}_{t},({\bm{z}}_{l},\emptyset),t)\big{)},$ (5)

### v2

# 1]FAIR at Meta 2]LTCI, Télécom Paris, IP Paris

@@ Abstract
Image codecs are typically optimized to trade-off bitrate vs. distortion metrics. At low bitrates, this leads to compression artefacts which are easily perceptible, even when training with perceptual or adversarial losses. To improve image quality and remove dependency on the bitrate we propose to decode with iterative diffusion models. We condition the decoding process on a vector-quantized image representation, as well as a global image description to provide additional context. We dub our model “PerCo” for “perceptual compression”, and compare it to state-of-the-art codecs at rates from 0.1 down to 0.003 bits per pixel. The latter rate is more than an order of magnitude smaller than those considered in most prior work, compressing a $512\!\times\!768$ Kodak image with less than 153 bytes. Despite this ultra-low bitrate, our approach maintains the ability to reconstruct realistic images. We find that our model leads to reconstructions with state-of-the-art visual quality as measured by FID and KID. As predicted by rate-distortion-perception theory, visual quality is less dependent on the bitrate than previous methods.
[…]
@@ 1 Introduction
Traditional image and video codecs are optimized for the rate-distortion function (Shannon, 1948), which minimizes the expected size of the data under a distortion constraint such as mean-squared error. Recent research has developed neural image compression methods that surpass handcrafted image compression codecs in terms of rate-distortion performance (Ballé et al., 2017, 2018; Cheng et al., 2020; He et al., 2022a). However, optimization for the rate-distortion function comes at a cost of “realism”, where realism is mathematically related to the statistical fidelity or $f$-divergence between the compressed image distribution and the true image distribution (Blau & Michaeli, 2018, 2019). The typical qualitative manifestation of unrealistic images is blurring.
@@
Generative modeling compensates for artefacts such as blurring by introducing a divergence term (Agustsson et al., 2019; Mentzer et al., 2020; Muckley et al., 2023), typically in the form of an adversarial discriminator loss, which improves human-perceptual performance (Mentzer et al., 2020). Such codecs are called “generative compression” codecs, and are evaluated in terms of the rate-distortion-realism tradeoff (Blau & Michaeli, 2019). A further result of rate-distortion-realism theory is the theoretical possibility of a perfect realism codec, i.e., a codec with no $f$-divergence, and zero FID across all rates, with no more than twofold increase in mean-squared error from the rate-distortion optimal codec. This result motivates research on perfect realism codecs, where original and reconstruction are different from each other, but it is not possible to tell which is which. Such codecs are particularly interesting for extremely low bitrate settings, where existing codecs introduce severe artefacts that are easily perceptible, see Figure 1. So far, the primary effort towards building such a codec was the work of Theis et al. (2022), who demonstrated lower FID scores than HiFiC (Mentzer et al., 2020) via a combination of a pretrained variational diffusion model and reverse channel coding. However, computational constraints prohibited application to images larger than $64\!\times\!64$ pixels.
@@
In this paper we make further progress towards perfect-realism codecs. Similar to Theis et al. (2022), we leverage a pretrained text-to-image diffusion model; however, rather than using the diffusion model for compression directly, we use it as a decoder in a VQ-VAE-like autoencoder. The (approximate) log-likelihood loss of diffusion models offers an alternative to the use of distortions such as MSE or LPIPS used to train most neural compression models that make strong assumptions on the conditonal distribution on images given the latents which are inappropriate at low bitrates, and lead to compression artefacts. For decoding, we sample from the conditional diffusion model, which allows us to obtain a set of reconstructions that reflect the uncertainty about the original source image. We augment the quantized image representation with a global image description that captures the high-level semantic information of the image. To this end, we use an automatically generated textual image description, e.g. using BLIP (Li et al., 2023). Alternatively, this can take the form of a global image feature extracted from a pre-trained image backbone. We demonstrate that this high-level information is helpful for compression performance at extremely low rates.
Our work is closely related to that of Lei et al. (2023), who explored text-conditioned diffusion decoders for image compression, and the only prior work that we are aware of that considers the same ultra-low bitrates as we do. In contrast, their work uses (automatically generated) binary contour maps as a spatial conditioning signal which carry little detail on colors and textures, and leads to reconstructions that are not faithful to the original. Experimentally we observe significantly improved FID and KID scores compared to competing methods. Moreover, we find that FID and KID are much more stable across bitrates than other methods, which aligns with our goal of image compression with perfect realism.
@@ 2 Related work
Neural image compression codecs. Ballé et al. (2017) demonstrated that an end-to-end neural image compression codec could outperform the classical JPEG codec in terms of rate-distortion performance. Ballé et al. (2018) enhance this approach by conditioning the latent on a “hyperprior” that encodes high-level information in the image, greatly improving performance. Follow-up works improved the latent conditioning mechanism (Minnen et al., 2018; Minnen & Singh, 2020; Cheng et al., 2020; He et al., 2022a). Optimization for rate-distortion alone leads to distributional mismatch between the compressed image distribution and the natural image distribution (Blau & Michaeli, 2018, 2019). Agustsson et al. (2019) and Mentzer et al. (2020) added an adversarial loss term and demonstrated its usefulness for human perception. Similarly He et al. (2022b) builds upon ELIC (He et al., 2022a) by adding perceptual losses including adversarial ones. Follow-up works have improved the discriminator architecture (Muckley et al., 2023), decode-time realism guidance (Agustsson et al., 2023), or investigated alternative autoencoders with discrete entropy models (El-Nouby et al., 2023). Xu et al. (2023) introduce the notion of conditional perceptual quality, where perceptual quality is dependent on side information. Our work is related to this, as the global textual or visual representation on which we condition the decoder can be seen as a form of side information.
@@
Diffusion models. Ho et al. (2020) improved the original diffusion probabilistic model (Sohl-Dickstein et al., 2015) that gave it impressive performance on image generation benchmarks. Dhariwal & Nichol (2021) demonstrated superior performance to GANs for image generation with classifier guidance, while Kingma et al. (2021) improved diffusion model likelihood estimation. Conditioning diffusion models on large text-encoder networks enables these models to generate realistic images from natural language prompts (Nichol et al., 2022; Saharia et al., 2022), and defining the diffusion process in an autoencoder latent space reduces the computational cost of the diffusion process. In our work we build upon large-scale pretrained text-conditioned diffusion models to benefit from the image generation capabilities embedded in them.
@@
Diffusion generative compressors. Ho et al. (2020) considered compression with diffusion models and reverse channel coding and the corresponding rate-distortion performance, but reported only experiments on $32\!\times\!32$ CIFAR-10 images and MSE distortion. Theis et al. (2022) demonstrated that the reverse channel coding approach gave FID numbers superior to HiFiC for $64\!\times\!64$ ImageNet images. Saharia et al. (2022) introduced a super-resolution diffusion model that removed artefacts from JPEG images. Ghouse et al. (2023) developed a complete version of a residual diffusion codec with images of high quality for a suite of codecs. Hoogeboom et al. (2023) showed similar results using a neural autoencoder baseline codec, and using a separate diffusion model that does not require any additional bits to refine the reconstruction. Yang & Mandt (2023) developed an alternative approach, where a hyperprior-based neural encoder was trained jointly with a diffusion decoder, also showing improved performance to HiFiC.
In Pan et al. (2022), they optimize a textual embedding on top of a pretrained text-to-image diffusion model. They also design a compression guidance method used at each denoising step to better reconstruct images. Lei et al. (2023) leverage text-conditioned diffusion models for image compression at very low bitrates, using prompt inversion to encode the image into text, and adding spatial detail via compressed binary contour maps. Our work is similar, but we replace computationally costly prompt inversion (Wen et al., 2023) with fast feed-forward image captioning (Li et al., 2023), and replace binary contour maps which carry little appearance information with end-to-end learned vector-quantized image features.
@@ 3.1 Overall framework
where $P_{\bm{x}}$ is the data distribution and $P_{{\bm{z}}|{\bm{x}}}$ is the posterior distribution of the quantized codes. The rate term $\mathcal{L}_{\textrm{R}}({\bm{z}})$
@@
estimates the bitrate either by means of scalar quantization and a continuous entropy model (Ballé et al., 2017), or using vector-quantization in combination with discrete entropy models (El-Nouby et al., 2023). For simplicity, in PerCo, we employ vector-quantization combined with a uniform entropy model leading to a constant rate controlled through two hyper-parameters: the number of elements that are quantized, and the codebook size. Therefore, the rate $\mathcal{L}_{\textrm{R}}$ can be considered fixed when training our models and can be dropped from the optimization formulation. Next, we introduce a new formulation of the distortion term, $\mathcal{L}_{\textrm{D}}$, enabling the integration of a pre-trained diffusion model, thereby achieving a higher level of perceptual quality. Our aim with this formulation is to leverage a large pre-trained diffusion model as a robust image prior, enabling realistic image reconstruction even at extremely low bitrates.
@@
Figure 2: Overview of PerCo. The LDM encoder maps an RGB image into the latent space of the diffusion model. The “hyper encoder” then maps the image to a hyper-latent with smaller spatial resolution, which is then vector-quantized and represented as a bitstream using uniform coding. The image captioning model generates a textual description of the input image, which is losslessly compressed, and processed by a text encoder to condition the diffusion model. The diffusion model reconstructs the input image in its latent space conditioned on the output of the text encoder and the hyper encoder. Finally, the LDM decoder maps the latent reconstruction back to RGB pixel space.
@@
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) approximate a data distribution by gradually denoising a unit Gaussian random variable. This is achieved by training a denoising function to reverse a diffusion process that takes an initial sample, denoted as ${\bm{x}}_{0}$, from the data distribution and adds small Gaussian noise over a large number of steps. In practice, a noise estimator $\epsilon_{\theta}({\bm{x}}_{t},t,{\bm{z}})$ implemented as a neural network is trained to denoise input data points at timesteps $t\in[0,T]$, where ${\bm{x}}_{t}\!=\!\sqrt{\alpha_{t}}{\bm{x}}_{0}\!+\!\sqrt{1-\alpha_{t}}\epsilon$ is a corrupted version of the original data point ${\bm{x}}_{0}$ with Gaussian noise $\epsilon$. The scalar $\alpha_{t}$ fixes the noise strength, varying from $\alpha_{0}\!\simeq\!1$ for no noise to $\alpha_{T}\!\simeq\!0$ for pure Gaussian noise. Finally, ${\bm{z}}$ is an optional condition, such as a text-prompt in text-to-image models. In our case, ${\bm{z}}$ is the quantized representation of the input image. Once the noise estimator is trained, the model can generate samples by drawing ${\bm{x}}_{T}\sim\mathcal{N}(0,I)$, and applying iterative denoising for $t\!=\!T,\dots,1$. To reduce the computational cost, we use a latent diffusion model (LDM) where the variables ${\bm{x}}_{t}$ are defined in the latent space of an autoencoder.
@@
$\mathcal{L}_{\textrm{Diff}}^{t}\propto\mathbb{E}_{P_{\bm{x}}}\mathbb{E}_{P_{{\bm{z}},{\bm{x}}_{t}|{\bm{x}}}}\mathop{\mathbb{E}}_{\epsilon\sim\mathcal{N}(0,1)}\lVert\bm{\epsilon}-\bm{\epsilon}_{\theta}({\bm{x}}_{t},{\bm{z}},t)\rVert_{2}^{2}.$ (3)
We obtain the final loss by taking the expectation of Equation 3 w.r.t. the steps $t$, and use ${\bm{v}}$ prediction objective of Salimans & Ho (2022), which is more stable during training.
The loss can be augmented with any image-level loss such as LPIPS (Zhang et al., 2018) to improve perceptual image similarity.
@@ 3.2 Encoding local and global context
Local spatial encoding. We encode the input image into a quantized tensor proceeding in three stages. First, we employ the LDM encoder $E(\cdot)$ to achieve a first dimension reduction. Assuming an input image of resolution $512\!\times\!512$, we obtain a feature map of dimension $4\!\times\!64\!\times\!64$. Second, we add a lightweight “hyper encoder” $H$ composed of several convolutional layers to project the LDM features to a lower spatial resolution tensor ${\bm{H}}_{s}$. The spatial resolution, $h\!\times\!w$, of ${\bm{H}}_{s}$ is chosen according to the target bitrate. Third, we proceed to quantization of ${\bm{H}}_{s}$ to obtain ${\bm{z}}_{l}$ via vector-quantization. As in VQ-VAE (van den Oord et al., 2017; Razavi et al., 2019), each of the $h\!\times\!w$ vectors ${\bm{h}}_{s}$ in ${\bm{H}}_{s}$ is mapped to an element in a codebook learned using the vector quantization loss:
@@
where $sg(.)$ is the stop-gradient operation, and ${\bm{z}}_{q}$ is the mapping of ${\bm{h}}_{s}$ to its closest codebook entry. For better codebook usage, we follow Yu et al. (2022) and set the output of the hyper encoder to be relatively low dimensional: 32 in our case. We also found it beneficial to replace the codebook loss by exponential moving average on the codebooks and thus the first term in Equation 4 is no longer used. The quantized encoder output is then upsampled to $64\!\times\!64$ resolution, to be channel-wise concatenated to the input of the noise estimation network of the diffusion model. The hyper-network is trained end-to-end while finetuning the diffusion model. For simplicity, we opt for uniform coding to map the quantized latents to a bitstream. Using $\log_{2}V$ bits to encode each element in the $h\!\times\!w$ sized hyper latent encoding, where $V$ is the codebook size, results in a total of $hw\log_{2}V$ bits.
@@
Global encoding with image captioning. While ${\bm{z}}_{l}$ can accurately encode local information, our experiments show that relying on ${\bm{z}}_{l}$ alone leads to unsatisfying realism. Therefore, we add a global encoding of the image ${\bm{z}}_{g}$ which provides additional context. For this purpose, we use an off-the-shelf state-of-the-art image captioning model, such as BLIP-2 (Li et al., 2023) or IDEFICS (Laurançon et al., 2023), which we keep fixed when training our model. Note that user-generated captions can also be used, as explored in our experiments. Similar to Lei et al. (2023), we losslessly compress the caption using Lempel-Ziv coding as implemented in the zlib library (zli, ) to obtain ${\bm{z}}_{g}$. Alternatively, we also explore the use of global image features extracted using an image backbone network.
@@
Decoding with a diffusion model. The quantized local image features in ${\bm{z}}_{l}$ are fed to the denoising U-Net of the diffusion model via concatenation with the latent features ${\bm{x}}_{t}$ of the current time step. To adapt to this change of dimension, we extend the kernel of the first convolutional layer of the U-Net by adding a number of channels that corresponds to the channel dimension of ${\bm{z}}_{l}$ which is randomly initialized. The global encoding ${\bm{z}}_{g}$ is losslessly decoded and passed to the diffusion model through cross-attention layers of the pre-trained diffusion model. Furthermore, we found it beneficial to use classifier-free guidance (Ho & Salimans, 2021), which we apply at inference time for the text conditioning ${\bm{z}}_{g}$, i.e. we contrast conditioning on local featues ${\bm{z}}_{l}$ alone vs. conditioning on both ${\bm{z}}_{l}$ and ${\bm{z}}_{g}$. This leads to the noise estimate:
$\hat{\epsilon}_{\theta}=\epsilon_{\theta}({\bm{x}}_{t},({\bm{z}}_{l},\emptyset),t)+\lambda_{s}\leavevmode\nobreak\ \big{(}\epsilon_{\theta}({\bm{x}}_{t},({\bm{z}}_{l},{\bm{z}}_{g}),t)-\epsilon_{\theta}({\bm{x}}_{t},({\bm{z}}_{l},\emptyset),t)\big{)},$ (5)